how do you simplify
x2+6x+5
______
x2-25
(x2= x squared or x to the second power)

Answer by fractalier(6550)

The key technique is to factor and cancel...so we do just that and get
(x^2 + 6x + 5) / (x^2 - 25) =
(x + 1)(x + 5) / (x - 5)(x + 5) =
(x + 1) / (x - 5)